# thanks: Corresponding author

### v1

@@ IV Numerical results
$\begin{split}&h_{1}=\frac{1}{2}\left\{\sigma^{x}\left(3\right)\sigma^{x}\left(2\right)+\sigma^{x}\left(2\right)\sigma^{x}\left(1\right)+\sigma^{x}\left(1\right)\sigma^{x}\left(0\right)\right.\\ &\left.+\sigma^{y}\left(3\right)\sigma^{y}\left(2\right)+\sigma^{y}\left(2\right)\sigma^{y}\left(1\right)+\sigma^{y}\left(1\right)\sigma^{y}\left(0\right)\right.\\ &\left.+\sigma^{x}\left(3\right)\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)\sigma^{x}\left(0\right)+\sigma^{y}\left(3\right)\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)\sigma^{y}\left(0\right)\right\},\\ &h_{2}=\frac{1}{2}\left\{\sigma^{z}\left(3\right)-\sigma^{z}\left(2\right)+\sigma^{z}\left(1\right)-\sigma^{z}\left(0\right)\right\},\\ &h_{3}=\left\{\sigma^{z}\left(3\right)+\sigma^{z}\left(2\right)+\sigma^{z}\left(1\right)+\sigma^{z}\left(0\right)\right\},\\ &+\frac{1}{4}\left\{\sigma^{z}\left(3\right)\sigma^{z}\left(2\right)+\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)+\sigma^{z}\left(1\right)\sigma^{z}\left(0\right)\right.\\ &\left.+\sigma^{z}\left(3\right)\sigma^{z}\left(0\right)\right\},\\ &h_{4}=\frac{1}{4}\left\{\sigma^{z}\left(3\right)\sigma^{z}\left(2\right)+\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)+\sigma^{z}\left(1\right)\sigma^{z}\left(0\right)\right.\\ &\left.+\sigma^{z}\left(3\right)\sigma^{z}\left(0\right)\right\}.\end{split}$ (26)
According to Eqs. (8) and (11), the observables in interest are,
$\begin{split}&\langle\bar{\psi}\psi\rangle=\langle h_{2}\rangle,\\ &\langle\bar{\psi}\gamma^{0}\psi\rangle=2+\frac{1}{2}\langle\sigma^{z}(0)+\sigma^{z}(1)+\sigma^{z}(2)+\sigma^{z}(3)\rangle.\\ \end{split}$ (27)
@@
In this work, the mass is fixed to $m=0.5a^{-1}$. For $H_{E}$, the range of $0.01a^{-1}\leq T\leq\;a^{-1}$ and $0\leq g^{2}\leq 3$ is considered, where $T=1/\beta$. For $H_{M}$, the range of $0.01a^{-1}\leq T\leq\;a^{-1}$ and $0\leq g^{2}\leq 5$ is considered. When $a^{-1}=100\;{\rm MeV}$, the parameters are $m=50\;{\rm MeV}$ and $1\;{\rm MeV}\leq T\leq 100\;{\rm MeV}$.
@@
Figure 1: $\langle\bar{\psi}\psi\rangle$ (the left panels) and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ (the right panels) as functions of $T$ and $g^{2}$ calculated by using exact diagonalization. The panels in the first row correspond to $H_{E}$, and the ones in the second row correspond to $H_{M}$, respectively.
@@
Since $N=4$, it is also possible to perform exact diagonalization to calculate $e^{-\beta H}$. To focus on the errors in QMETTS, the discritization errors are ignored. In other words, the observables $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ are calculated using the Hamiltonian in Eq. (25), so that the deviations between the QMETTS and the exact diagonalization is purely due to the errors in QMETTS. The result of $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ are shown in Fig. 1. It can be seen that, the chiral symmetry is broken at either high temperatures or weak couplings. At low temperatures, the dependence of chiral condensation on the coupling constant is manifest in the form of distinct plateaus, which correspond to different topological charges. The values of the topological charge are integers at low temperatures, therefore exhibiting a topological structure.
@@
The goal of our study is to reproduce the above results using QMETTS. Since $N=4$ is small, we use a complete set of bases of the Hilbert space, $\Phi_{i}=|i\rangle$, for $0\leq i\leq 15$. The range of temperature is $0.1a^{-1}\leq T\leq 2a^{-1}$, in other words, $0.5\leq\beta\leq 10$ with $\Delta\beta=0.25$ and $K=20$, where $K=\beta_{max}/(2\Delta\beta)$ is defined in previous section. Therefore, the temperatures in consideration is $a^{-1}/\left(2k\Delta\beta\right)$ with $1\leq k\leq K$. The number of steps in Trotter decomposition is $t=10$. In the evaluation, only the $\hat{\sigma}_{j}$ terms in Eq. (17) with $|a_{j}|>0.001$ are kept. In the case of $H_{M}$, the values of the coupling constant are chosen as $g^{2}=0,0.2,0.4,\ldots,5.0$. In the case of $H_{E}$, the values of the coupling constant are chosen as $g^{2}=0,0.1,0.2,\ldots,3.0$.
Figure 2: $\langle\bar{\psi}\psi\rangle$ (the left panels) and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ (the right panels) as functions of $T$ and $g^{2}$ calculated by using QMETTS. The panels in the first row correspond to $H_{E}$, and the ones in the second row correspond to $H_{M}$, respectively.
@@
The circuit is implemented using qiskit [74], and the results are obtained by using a simulator. The observables $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ at different temperatures and coupling constants are shown in Fig. 2. For simplicity, the parameters to calculate $a_{j}$ and $C(\beta)$ used in the evaluation are measured exactly which is only possible using a simulator, while $O_{i}$ in Eq. (24) are measured for $1024$ times. As can be seen from Figs. 1 and 2, the outcomes yielded by QMETTS are largely in alignment with those obtained through exact diagonalization. Among these results, the breaking and restoration of chiral symmetry, the plateaus in chiral condensation, and the behavior of topological charges are all consistent. In QMETTS, the jitter in the values is primarily attributable to the number of measurements. The edges of the plateaus at low temperatures are observed to be less sharp than in exact diagonalization, due to the fact that in QMETTS we only evolve up to $\beta=10$, in contrast to the exact diagonalization where $\beta=100$.
Figure 3: $\langle\bar{\psi}\psi\rangle$ (the left panels) and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ (the right panels) as functions of $g^{2}$ at $T=0.1a^{-1}$. The solid curves are the results of exact diagonalization, while the dots are the results obtained by the QMETTS. The panels in the first row correspond to $H_{E}$, and the ones in the second row correspond to $H_{M}$, respectively.
In the QMETTS approach, the behaviors at large $\beta$ suffer from the accumulated errors from each $k\Delta\beta$ evolution. Whether the interesting behaviors at low temperatures can be reproduced is investigated. At $T=0.1a^{-1}$ (which is $T=10\;({\rm MeV})$), the observables are compared in Fig. 3. It can be seen that, even when $\beta=10$, the results from QMETTS are consistent with the results from exact diagonalization well.
@@
V Summary

### v2

@@ 4 Numerical results
$\begin{split}h_{1}&=\frac{1}{2}\left\{\sigma^{x}\left(3\right)\sigma^{x}\left(2\right)+\sigma^{x}\left(2\right)\sigma^{x}\left(1\right)+\sigma^{x}\left(1\right)\sigma^{x}\left(0\right)\right.\\ &\left.+\sigma^{y}\left(3\right)\sigma^{y}\left(2\right)+\sigma^{y}\left(2\right)\sigma^{y}\left(1\right)+\sigma^{y}\left(1\right)\sigma^{y}\left(0\right)\right.\\ &\left.+\sigma^{x}\left(3\right)\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)\sigma^{x}\left(0\right)+\sigma^{y}\left(3\right)\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)\sigma^{y}\left(0\right)\right\},\\ h_{2}&=\frac{1}{2}\left\{\sigma^{z}\left(3\right)-\sigma^{z}\left(2\right)+\sigma^{z}\left(1\right)-\sigma^{z}\left(0\right)\right\},\\ h_{3}&=\left\{\sigma^{z}\left(3\right)+\sigma^{z}\left(2\right)+\sigma^{z}\left(1\right)+\sigma^{z}\left(0\right)\right\}\\ &+\frac{1}{4}\left\{\sigma^{z}\left(3\right)\sigma^{z}\left(2\right)+\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)+\sigma^{z}\left(1\right)\sigma^{z}\left(0\right)+\sigma^{z}\left(3\right)\sigma^{z}\left(0\right)\right\},\\ h_{4}&=\frac{1}{4}\left\{\sigma^{z}\left(3\right)\sigma^{z}\left(2\right)+\sigma^{z}\left(2\right)\sigma^{z}\left(1\right)+\sigma^{z}\left(1\right)\sigma^{z}\left(0\right)+\sigma^{z}\left(3\right)\sigma^{z}\left(0\right)\right\}.\end{split}$ (22)
According to Eq. (7), the observables of interest are,
[…]
@@
In this work, the mass is fixed at $m=0.5a^{-1}$. For $H_{E}$, the range of $0.01a^{-1}\leq T\leq\;a^{-1}$ and $0\leq g^{2}\leq 3$ is considered, where $T=1/\beta$. For $H_{M}$, the range of $0.01a^{-1}\leq T\leq\;a^{-1}$ and $0\leq g^{2}\leq 5$ is considered. When $a^{-1}=100\;{\rm MeV}$, the parameters are $m=50\;{\rm MeV}$ and $1\;{\rm MeV}\leq T\leq 100\;{\rm MeV}$.
@@
Figure 1: Chiral condensation $\langle\bar{\psi}\psi\rangle$ (the left panels) and fermion number (topological charge) $\langle\bar{\psi}\gamma^{0}\psi\rangle$ (the right panels) as functions of $T$ and $g^{2}$ calculated by using exact diagonalization. The panels in the first row correspond to $H_{E}$, and the ones in the second row correspond to $H_{M}$, respectively. Chiral symmetry breaks at low temperatures or weak coupling strengths. At low temperatures, the topological charge takes integer values, indicating a topological structure.
@@
Since $N=4$, exact diagonalization can be used to compute $e^{-\beta H}$. To focus on QMETTS numerical results, discussion of discretization errors is deferred to the next section. The observables $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ are calculated using the Hamiltonian in Eq. (21). The results of $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ using exact diagonalization are shown in Fig. 1. It can be seen that, the chiral symmetry is broken at either low temperatures or weak couplings. At low temperatures, the dependence of chiral condensation on the coupling constant is manifest in the form of distinct plateaus, which correspond to different topological charges. The values of the topological charge are integers at low temperatures, therefore exhibiting a topological structure.
@@
The goal of our study is to reproduce the above results using QMETTS. Since $N=4$ is small, we use a complete set of bases of the Hilbert space, $\Phi_{i}=|i\rangle$, for $0\leq i\leq 15$. The range of temperature is $0.1a^{-1}\leq T\leq 2a^{-1}$, in other words, $0.5\leq a^{-1}\beta\leq 10$ with $a^{-1}\Delta\beta=0.25$ and $K=20$, where $K=\beta_{max}/(2\Delta\beta)$ is defined in previous section. Therefore, the temperatures in consideration is $1/\left(2k\Delta\beta\right)$ with $1\leq k\leq K$. The number of steps in Trotter decomposition is $t=10$. In the evaluation, only the $\hat{\sigma}_{j}$ terms in Eq. (13) with $|c_{j}|>0.001$ are kept. In the case of $H_{M}$, the values of the coupling constant are chosen as $g^{2}=0,0.2,0.4,\ldots,5.0$. In the case of $H_{E}$, the values of the coupling constant are chosen as $g^{2}=0,0.1,0.2,\ldots,3.0$.
Figure 2: $\langle\bar{\psi}\psi\rangle$ (the left panels) and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ (the right panels) as functions of $T$ and $g^{2}$ calculated by using QMETTS. The panels in the first row correspond to $H_{E}$, and the ones in the second row correspond to $H_{M}$, respectively. For clarity, statistical uncertainties are omitted , with their analysis deferred to the following section. Comparison with Fig. 1 demonstrates that QMETTS generally captures the evolution of both the chiral condensation and topological charge with varying temperature and interaction strength.
@@
The circuit is implemented using qiskit qiskit2024 , and the results are obtained by using a simulator on a classical computer. The observables $\langle\bar{\psi}\psi\rangle$ and $\langle\bar{\psi}\gamma^{0}\psi\rangle$ at different temperatures and coupling constants are shown in Fig. 2. For simplicity, the parameters to calculate $c_{j}$ and $C(\beta)$ used in the evaluation are measured exactly which is only possible using a simulator, while $O_{i}$ in Eq. (20) are measured for $r=1024$ times. As can be seen from Figs. 1 and 2, the outcomes yielded by QMETTS are largely in alignment with those obtained through exact diagonalization. Among these results, the breaking and restoration of chiral symmetry, the plateaus in chiral condensation, and the behavior of topological charges are all consistent. In QMETTS, the jitter in the values is primarily attributable to the number of measurements. The edges of the plateaus at low temperatures are observed to be less sharp than in exact diagonalization, due to the fact that in QMETTS we only evolve up to $a^{-1}\beta=10$, in contrast to the exact diagonalization where $a^{-1}\beta=100$.
@@
5 A discussion on the errors
There are three primary sources of errors in quantum simulations of the Thirring model, the model-related errors, algorithmic errors, and implementation errors. Due to the substantial circuit depth required for low-temperature simulations, which currently prevents physical implementation, this section focuses on model-related and algorithmic errors. In this work, we neglect noise effects to demonstrate that the QITE algorithm can inherently handle low-temperature systems with topological phase transitions under ideal conditions. The study of noise resilience is deferred to future research.
5.1 The systematic errors of the model
By applying Jordan-Wigner transformations, the Hamiltonians of the Thirring model are mapped to those in Eq. (21). However, Hamiltonians in Eq. (21) are subject to both discretization errors and finite-volume effects.
Figure 3: At $a^{-1}\beta=1$ ($T=100\;{\rm MeV}$) and different $g^{2}$, the average chiral condensate $\langle\bar{\psi}\psi\rangle/N$ and topological charge $\langle\bar{\psi}\gamma_{0}\psi\rangle/N$ as functions of $1/V$ for lattice sizes $N=4$, 66, 88, $10$, and $12$, along with corresponding linear extrapolations. The upper panels are the results of $H_{M}$, the bottom panels are results of $H_{E}$, respectively. It can be observed that finite-volume effects are negligible at higher temperatures.
Figure 4: Same as Fig. 3 but for $a^{-1}\beta=10$ ($T=10\;{\rm MeV}$). The effect of finite volume is significant at smaller $g^{2}$, especially for the case of $H_{M}$ (the upper panels). However, the regime with $1/V\to 0$ can be effectively captured through linear extrapolations.
To analyze finite volume effects, we examine the asymptotic behavior of observables via exact diagonalization across increasing system sizes. In order to consider only the influence of the finite volume, we calculate the average of chiral condensation and the average of topological charge, i.e., $\langle\bar{\psi}\psi\rangle/N$ and $\langle\bar{\psi}\gamma_{0}\psi\rangle/N$. The cases of $N=4$, 66, 88, $10$ and $12$ are studied, the results at $a^{-1}\beta=1$ and $10$ for $g^{2}=1$, $2$ and $3$ are shown in Figs. 3 and 4.
The results are shown as functions of $1/V$ where $V=aN$ is the volume. It can be found that, at higher temperatures ($a^{-1}\beta=1$), the effect of finite volume is minor, and even results at $N=4$ closely match those at $N=12$. In contrast, at lower temperatures ($a^{-1}\beta=10$), the effect of finite volume becomes significant for smaller $g^{2}$, particularly for the case of $H_{M}$. This is understandable because when topological order arises, the larger the volume, the more difficult it is for the system to switch between different topological sectors. When $g^{2}$ is relatively large, the topological sector of the final state is closer to the initial state, whereas when $g^{2}$ is relatively small, the transition from the initial state to the final state requires switching between different topological sectors, explaining the observed sensitivity. In Figs. 3 and 4, we simultaneously present results from linear extrapolation. It can be observed that although the finite volume effect has significant impact, the worst case scenario is the chiral condensate at $g^{2}=1$ and $a^{-1}\beta=10$, where for $N=4$ $\langle\bar{\psi}\psi\rangle/N\approx-0.07$ and for $N=12$ one has $\langle\bar{\psi}\psi\rangle/N\approx-0.13$, the regime with $1/V\to 0$ can be effectively captured through linear extrapolations.
Figure 5: .At $T=100\;{\rm MeV}$ and different $g^{2}$, the average chiral condensate $\langle\bar{\psi}\psi\rangle/N$ and topological charge $\langle\bar{\psi}\gamma_{0}\psi\rangle/N$ as functions of $a$ for lattice sizes $N=4$, 66, 88, $10$, and $12$, along with corresponding linear extrapolations. The upper panels are the results of $H_{M}$, the bottom panels are results of $H_{E}$, respectively. The convergence behavior of observables is not purely linear, therefore the linear extrapolation is performed with the data points of $N=8,10,12$.
Figure 6: .Same as Fig. 5 but for $T=10\;{\rm MeV}$. For chiral condensation, the slope of the linear extrapolation is smaller than those at $T=100\;{\rm MeV}$. For the $H_{E}$ (the bottom panels), convergence exhibits oscillatory behavior at certain $g^{2}$ values. To obtain more reliable results, additional data points are required for linear extrapolation.
Similarly the discretization error is also studied by using $N=4$, 66, 88, $10$ and $12$ lattices but with $a^{-1}=100$, $150$, $200$, $250$ and $300\;{\rm MeV}$. To make sure the model is the same one under different lattice spacings, we consider the cases of $T=100\;{\rm MeV}$ and $10\;{\rm MeV}$ with different $\beta$. For example, for $N=4$, the two cases correspond to $a^{-1}\beta=1$ and $10$, for $N=6$, the two cases correspond to $a^{-1}\beta=1.5$ and $15$, etc. Apart from that, $m=50\;{\rm MeV}$ is fixed and $am$ is different for different lattice extents. The average of chiral condensation and the average of topological charge at $T=100\;{\rm MeV}$ and $10\;{\rm MeV}$ are shown in Figs. 5 and 6, respectively.
Compared with the finite volume effect, the chiral condensation convergent faster at lower temperatures. Since the finite volume effect can be viewed as inferred cut off, while the discretization error can be viewed as the ultra-violate cut off, this result makes intuitive sense. Furthermore, as the lattice spacing decreases, the convergence behavior of observables is not purely linear, exhibiting an additional exponential convergence or oscillatory convergence. Given that higher-order contributions become negligible for sufficiently small $a$, we perform the extrapolation exclusively on the $N=8,10,12$ data points while retaining the linear extrapolation. The results of the linear extrapolations are also shown in Figs. 5 and 6. The results indicate that linear extrapolation yields reasonable outcomes in most cases, with the exception of scenarios dominated by oscillatory or exponential convergence, where it overestimates discretization errors. It can be expected that the extrapolation results become more reliable as long as a few number of additional data points are incorporated.
5.2 The systematic errors of the method
When using QMETTS, the system errors arise from approximating a thermal state with $|\phi_{i}\rangle/\sqrt{C_{i}}$, omitting the contribution of the $A$ matrix with coefficients $|c_{j}|<0.001$, and the Trotter expansion. These deviations can be analyzed numerically by directly comparing the results of quantum simulation with exact diagonalization. In the following analysis, statistical errors are also considered. For both chiral condensation and fermion number, the statistical error of measuring the observables by repeating the circuit by $r$ times can be estimated as,
$\begin{split}&\epsilon_{st.}=\frac{1}{2}\frac{\sqrt{\sum_{i=1}^{16}\sum_{n=0}^{3}\left(\frac{C_{\phi_{i}}}{\sum_{j=1}^{16}C_{\phi_{j}}}\right)^{2}\left(1-\left\langle\phi_{i}\left(\frac{\beta}{2}\right)\right|\sigma_{z}(n)\left|\phi_{i}\left(\frac{\beta}{2}\right)\right\rangle^{2}\right)}}{\sqrt{r}}.\end{split}$ (24)
Figure 7: .Comparison of QMETTS-derived observables with exact diagonalization results across different $g^{2}$ values as $\beta$ increases. Error bars represent statistical uncertainties from measurements estimated using Eq. (24). The upper panels are the results of $H_{M}$, the bottom panels are results of $H_{E}$, respectively. With noise ignored, the deviation between QMETTS and exact diagonalization does not accumulate with the evolution of $\beta$.
For different $g^{2}$, the results of the observables measured with statistic errors by using QMETTS as well as the results from exact diagonalization along the evolution of $\beta$ are shown in Fig. 7. It should be emphasized that the circuit is simulated on a classical computer, and the effect of noise is not considered. It can be shown that in this case, the deviation between QMETTS and exact diagonalization does not accumulate with growing $\beta$. This can be understood by one of the inherent advantages of QITE-based algorithms, that process of imaginary time evolution is akin to “cooling” a quantum system to its lowest energy configuration such that an initial state with any non-negligible overlap with the true ground state is guaranteed to evolve towards that ground state Kolotouros:2024inn . This provides an advantage over other quantum-classical hybrid algorithms using classical optimizers, which can become trapped in local minima during optimization.
Figure 8: Same as Fig. 7 but for evaluating to a same $\beta$ with growing coupling constant $g^{2}$. It is found that, the deviation does not accumulate with growing $g^{2}$ either. Note that, at $a^{-1}\beta=10$ ($T=10\;{\rm MeV}$), the results are consistent with the exact diagonalization, indicating that the QMETTS can reproduce the behavior of massive Thirring model at low temperatures.
Observables computed via QMETTS and exact diagonalization are also compared across different $g^{2}$, as shown in Fig. 8. It can be seen that, the deviation does not accumulate with growing $g^{2}$ either. The reason will be explored latter. Within the range of coupling strengths $g^{2}$ considered, QITE proves applicable as a non-perturbative method. The results at $a^{-1}\beta=10$ ($T=10\;{\rm MeV}$) are consistent with the exact diagonalization, demonstrating the capabilities of QMETTS to capture the low temperature physics.
Figure 9: Distributions of $\Delta\langle O\rangle$ for all measurements. The upper panels are the results of $H_{M}$, the bottom panels are results of $H_{E}$, respectively. The deviations distributed within the range $\left|\Delta\langle O\rangle\right|<0.06$, and are approximately centered at zero.
In the case of $H_{M}$, with $K=20$ and $g^{2}=0,0.2,\ldots,5.0$, there are $20\times 26$ data points measured for each observable, similarly, in the case of $H_{E}$, $g^{2}=0,0.1,\ldots,3.0$, there are $20\times 31$ data points. Defining $\Delta\langle O\rangle=\langle O\rangle_{\rm QMETTS}-\langle O\rangle_{\rm exact}$ where $\langle O\rangle_{\rm QMETTS}$ and $\langle O\rangle_{\rm exact}$ are values of observables obtained by using QMETTS and exact diagonalization, respectively, the distributions of $\Delta\langle O\rangle$ over the above mentioned data points are shown in Fig. 9. It can be seen that, the deviations distributed within the range $\left|\Delta\langle O\rangle\right|<0.06$, and are approximately centered at zero. The characteristic pattern of error distribution validates the reliability of QMETTS.
Figure 10: The number of required Pauli tensor products during $\beta$ evolution. Given the variation in Pauli tensor counts depending on both $g^{2}$ and initial states, only the scenarios with the maximum required count are shown. The circuit depths remain substantial when evolving to larger $\beta$ values.
Another source of systematic error and computational cost not addressed is the calculation of the coefficients $c_{j}$. This part of computational cost increases with the size of the system. Since the operator $A$ is a linear combination of tensor products of Pauli matrices, the circuit depth can be roughly estimated by their number (note that actual depth varies for different tensor products and depends on physical qubit topology, these variations are neglected). For different initial states $|\phi_{i}\rangle$ and different $g^{2}$, the required number of Pauli tensor products differs. We consider the maximum number of Pauli tensor products required when evolving to $a^{-1}\beta=10$ across all initial states, which is denoted as $n_{A}$. $n_{A}$ serves as a rough estimate of both circuit depth and measurement complexity. For $H_{M}$ and $H_{E}$, the $n_{A}$ throughout $\beta$ evolution are depicted in Fig. 10. It can be seen that the circuit depths required for evolution to larger $\beta$ values remain beyond the capabilities of current quantum hardware.
Since the deviation between QMETTS and exact diagonalization arises from thermal state approximations and Trotter errors, the deviation can be estimated to correlate with $n_{A}$. We find that to evolute to $a^{-1}\beta=10$, the values of $n_{A}$ remain comparable across different $g^{2}$. For $g^{2}=0$, $n_{A}=764$ are required. In the case of $H_{M}$ at $g^{2}=5$, the requirement is $n_{A}=774$. In the case of $H_{E}$ at $g^{2}=3$, up to $n_{A}=892$ terms are needed. Consequently, this explains why the deviation does not accumulate with growing $g^{2}$ as shown in Fig. 8.
If we drop the periodic boundary conditions, for small $\beta$’s, the calculations of the coefficients can be simplified. Because at this point, the $H_{M}$ becomes an $XX$ model with an external transverse field in $1+1$ dimension, and $H_{E}$ becomes an $XXZ$ model with an external longitudinal field in $1+1$ dimension. Each term in the Hamiltonian acts on at most 2 qubits (2-local), the $A$ operator only affects the domain around these two qubits been acted by the term Motta:2019yya . However this simplification disappears with the long-range correlation of the wave function with the increase of $\beta$. In this case, one can truncate the $A$ matrix using schemes such as inexact-QITE Motta:2019yya or extended local approximation or non-local approximation Nishi:2021xhf . The truncation of the $A$ operator brings not only a simplification of the calculation of coefficients, but also a reduction of the circuit depth, as a result the computational complexity grows with the size of the correlation domain (or the order of tensor products in $A$ operator in the case of inexact-QITE) instead of the size of the physics system to be simulated Motta:2019yya ; Nishi:2021xhf . Given our exploration into the low-temperature regime, where circuit depth becomes substantial, in our case implementation remains infeasible on real NISQ devices. To address the issue of excessive circuit depth, several algorithmic approaches have been developed, for example by reincorporating variational machinery into the QITE algorithm, as exemplified by Refs. Kumar:2024kij ; Gomes:2021ckn . Improvements to QITE-algorithms are still active areas Silva:2021jaj ; Hejazi:2023fiq ; Gluza:2024lqq , and the massive thirring model can provide a test ground on those improvements.
6 Summary